### v3

# Error Correction Code Transformer: From Non-Unified to Unified

Abstract
Channel coding is vital for reliable data transmission in modern wireless systems, and its significance will increase with the emergence of sixth-generation (6G) networks, which will need to support various error correction codes. However, traditional decoders were typically designed as fixed hardware circuits tailored to specific decoding algorithms, leading to inefficiencies and limited flexibility. To address these challenges, this paper proposes a unified, code-agnostic Transformer-based decoding architecture capable of handling multiple linear block codes, including Polar, Low-Density Parity-Check (LDPC), and Bose-Chaudhuri-Hocquenghem (BCH), within a single framework. To achieve this, standardized units are employed to harmonize parameters across different code types, while the redesigned unified attention module compresses the structural information of various codewords. Additionally, a sparse mask, derived from the sparsity of the parity-check matrix, is introduced to enhance the model’s ability to capture inherent constraints between information and parity-check bits, resulting in improved decoding accuracy and robustness. Extensive experimental results demonstrate that the proposed unified Transformer-based decoder not only outperforms existing methods but also provides a flexible, efficient, and high-performance solution for next-generation wireless communication systems.
1 Introduction
With the continuous evolution of wireless communication technologies, the imminent arrival of sixth-generation (6G) wireless communications heralds a paradigm shift towards ultra-reliable, low-latency communications and notably enhanced data rates  [1, 2, 3]. To fulfill the ambitious performance requirements of 6G networks, probabilistic coding schemes are anticipated to play a crucial role. In particular, linear block codes such as Polar codes  [4] and Low-Density Parity-Check (LDPC) codes  [5] are expected to maintain their competitive edge established in fifth-generation (5G) wireless communications. These codes have also emerged as promising candidates for efficient error correction in 6G  [6]. Additionally, another algebraic linear block code, known as the Bose-Chaudhuri-Hocquenghem (BCH) code  [7], is often used as an outer code to enhance decoding performance.
1.1 Prior Works
Looking at 5G and earlier wireless communication systems, linear block code decoders were typically designed as fixed hardware circuits tailored to specific decoding algorithms. Looking ahead to 6G, this trend is likely to persist  [8]. For instance, the Successive Cancellation (SC) decoding algorithm  [4] for Polar codes uses a sequential method that traverses a decision tree to obtain the decoding result. This approach simplifies the hardware implementation of Polar codes by pruning binary trees  [9, 10, 11]. In contrast, the Sum-Product (SP) decoding algorithm  [12] for LDPC codes uses iterative message passing, allowing for parallel data processing. This fundamentally differs from the sequential method used for Polar codes in hardware implementation  [13]. Furthermore, BCH codes typically use algebraic decoding techniques  [14], classified as hard decoding, which differs greatly from soft decoding methods like SC and SP. Overall, these algorithms provide strong error correction capabilities. However, due to the non-unified nature of the sequential and parallel architectures of Polar and LDPC codes, and the gap between hard and soft decoding in BCH codes, unique hardware circuits must be customized and fine-tuned for each specific linear block code. This may lead to significant engineering effort and inefficient use of hardware resources.
In an effort to reduce hardware resource consumption and establish a quasi-unified decoding framework, three methods have been explored: resource sharing, Ordered-Statistics Decoding (OSD), and Artificial Intelligence (AI)-driven approaches. Resource sharing involves time-sharing storage and computing units. Ordered-Statistics Decoding leverages the reliability ordering of received symbols to efficiently decode various linear block codes by systematically evaluating the most likely codewords. Meanwhile, AI-driven methods offer a convenient way to unfold traditional algorithms like Belief Propagation (BP).
First, resource-sharing methods, such as those in  [15] and  [16], endeavor to integrate decoders for Polar and LDPC codes into a shared architecture based on traditional decoding algorithms. Both approaches aim to conserve hardware resources by temporally multiplexing the computational and storage units shared by the decoder through reconfiguration. However, due to the distinct decoding algorithms for each code type, achieving a truly unified architecture with such decoders is generally challenging.
Second, ordered-statistics decoding methods, initially explored in  [17, 18, 19], provide a near-Maximum Likelihood (ML) decoding approach for linear block codes by exploiting the reliability ordering of received symbols. The algorithm operates by sorting the received signal positions based on their reliability, permuting the code’s generator matrix accordingly, and systematically evaluating the most likely codewords through a limited search. OSD delivers near-ML performance with significantly less complexity than exhaustive ML decoding, making it well-suited for short to moderate-length codes. However, its drawback lies in the computational complexity, which grows exponentially with the search order (e.g., order-$l$). For instance, an order-4 OSD may require evaluating thousands of test codewords, potentially leading to increased latency and resource demands.
Third, in recent years, AI-driven decoders have gained attention as an alternative. While model-based AI methods leverage traditional decoder algorithms and rely on neural networks to simulate belief propagation, their performance improvements are often limited by the constraints of conventional models  [20],  [21]. In contrast, model-free approaches offer a more flexible and scalable solution. Unlike model-based methods, which depend on predefined mathematical models, model-free approaches assume no prior knowledge of the encoder or decoder. Instead, they rely on large amounts of labeled data to learn the decoding process directly, bypassing the limitations of conventional algorithms  [22].
For instance, inspired by the success of Transformers in Natural Language Processing (NLP)  [23], a model-free neural decoder called the Error Correction Code Transformer (ECCT) was proposed  [24]. ECCT employs the Transformer’s self-attention mechanism, where multiple attention heads operate in parallel to process sequential data and capture long-range dependencies, effectively linking input Log-Likelihood Ratios (LLRs) to the decoder’s output. The inherent connection between Transformers and decoders lies in their shared ability to translate sequences—whether from one language to another or from LLRs back to the original information bits. However, ECCT’s flexibility comes at a cost, requiring frequent parameter adjustments for different code types, lengths, and rates, which increases training complexity and hardware overhead, hindering a unified decoding architecture. Building on this,  [25] introduced a Foundation Model for Error Correction Codes (FECCT) to improve ECCT’s generalization to unseen codewords. Yet, FECCT currently faces the challenge of high computational complexity in its self-attention mechanism, further complicating its practical implementation.
1.2 Contributions
Thus, to reduce engineering efforts, training costs, and hardware consumption, we propose a unified, code-agnostic Transformer-based decoding architecture capable of handling multiple linear block codes, including Polar, LDPC, and BCH, within a single framework. To achieve this, we designed new standardized units and a unified attention module. Additionally, a sparse masking mechanism was developed to enhance decoding performance and stability  [^1] Simulation codes will be provided to reproduce the results in this paper: http://oa.ee.tsinghua.edu.cn/dailinglong/publications/publications.html. Specifically, the main contributions of this paper are summarized as follows.
  1. 1.
    Unified Decoding Architecture: We introduce a unified, code-agnostic Transformer decoding architecture, capable of seamlessly integrating various linear block codes. This architecture features standardized units to align various code lengths and, for the first time, designs a unified attention module that learns to compress the structural information of all codewords, thereby achieving a code-agnostic decoder. This architecture obviates the necessity for distinct hardware circuits for different decoding algorithms, reducing engineering efforts and hardware resource inefficiency.
  2. 2.
    Standardized Units: Standardized units are used to align parameters such as code length and code rate across different code types. Specifically, after receiving the channel output, we pad different codewords and their corresponding syndromes with zeros up to the maximum length. Through this unit, the Transformer neural decoder can accommodate any codeword within the maximum length, allowing for mixed training in a unified manner, thereby reducing training overhead.
  3. 3.
    Unified Attention: We redesigned the self-attention module to create a unified attention mechanism that allows different attention heads to share attention scores. This approach contrasts with traditional self-attention mechanisms, where each head independently computes its scores using separate learned weights. The unified attention compresses the structural information of different codewords through learning and shares parameters across various heads in the Transformer, resulting in a more cohesive decoding architecture.
  4. 4.
    Performance Enhancement: By introducing a sparse mask, which leverages the sparsity of the parity-check matrix, we have significantly enhanced the model’s ability to capture inherent constraints between information and parity-check bits. This has led to a substantial improvement in decoding performance and the achievement of a state-of-the-art design.
  5. 5.
    High Computational Efficiency: The unified attention module in our proposed Transformer-based decoding architecture demonstrates linear computational complexity by eliminating the need to project the input into higher dimensions. This results in a significant reduction in complexity from the traditional $\mathcal{O}(N^{2})$ to a more efficient $\mathcal{O}(N)$, where $N$ represents the sequence length. Additionally, by employing our sparse mask, we further achieve an $86\%$ reduction in computational complexity. This efficiency gain enables faster processing and lower power consumption, making the architecture suitable for real-time applications in next-generation wireless communication systems.
The rest of this paper is organized as follows. Section II provides background knowledge, including the vanilla Transformers, channel encoders, as well as pre- and post-processing involved to address the issue of overfitting. Section III details our proposed unified code-agnostic Transformer. Section IV presents the results of training and inference. Section V analyzes the impact of the proposed optimization algorithms. Finally, Section VI concludes.
Notations: Bold lowercase and uppercase letters denote vectors (e.g., $\mathbf{v}$) and matrices (e.g., $\mathbf{M}$), respectively, while scalars use regular letters (e.g., $x$). $\mathbb{F}_{2}$ represents the Galois field of order 2, and $\mathbb{R}$, $\mathbb{N}$ denote the real and natural numbers. The transpose is indicated by $\cdot]^{T}$, and $\mathcal{N}(\mu,\sigma^{2})$ denotes a Gaussian distribution with mean $\mu$ and variance $\sigma^{2}$. The function $\text{bipolar}(x)=1-2x$ maps $x\in{0,1}$ to bipolar values 1,1{1,-1}, with its inverse $\text{bin}(x)=\frac{1-x}{2}$ mapping 1,1{1,-1} back to 0,1{0,1}. The operator $\text{size}(\cdot)$ denotes the number of elements in a vector or matrix, and $\text{sign}(\cdot)$ represents the sign function, defined as $\text{sign}(x)=1$ if $x>0$, 0 if $x=0$, and 1-1 if $x<0$. For a linear block code with code length $n$ and information bit length $k$, we denote it as $(n,k)$.
2 Background
For a better understanding of the proposed unified error correction code Transformer, this section will provide background knowledge on vanilla Transformers and Forward Error Correction (FEC) codes.
2.1 Vanilla Transformers
The vanilla Transformer, a sequence-to-sequence model, comprises an encoder and decoder, each with $L$ identical blocks. Each block integrates multi-head self-attention (MHA), a position-wise feed-forward network (FFN), residual connections  [26] to mitigate gradient issues, and layer normalization  [27] for stability and faster convergence, as shown in Figure 1.
Figure 1: Single-layer vanilla Transformer encoder architecture.
The process begins with embedding a one-dimensional input signal $\textbf{{x}}\in\mathbb{R}^{N}$ into a higher-dimensional space:
$\textbf{{X}}_{e}=\text{Embedding}(\textbf{{x}})=\text{diag}(\textbf{{x}})\cdot\textbf{{W}}$ (1)
where $\text{diag}(\textbf{{x}})\in\mathbb{R}^{N\times N}$ is a diagonal matrix with x on the diagonal, and $\textbf{{W}}\in\mathbb{R}^{N\times d_{k}}$ is a trainable weight matrix, producing $\textbf{{X}}_{e}\in\mathbb{R}^{N\times d_{k}}$ with $d_{k}$-dimensional embeddings for each position.
This embedding feeds into the MHA mechanism, which correlates sequence elements through:
$\text{Attn}\left(\textbf{{Q}},\textbf{{K}},\textbf{{V}}\right)=\text{Softmax}\left(\frac{\textbf{{Q}}\cdot\textbf{{K}}^{T}}{\sqrt{d_{k}}}\right)\cdot\textbf{{V}}$ (2)
extended to multiple heads as:
$\text{MHA}\left(\textbf{{Q}},\textbf{{K}},\textbf{{V}}\right)=\text{Concat}\left(\text{head}_{1},\cdots,\text{head}_{H}\right)\cdot\textbf{{W}}^{\textit{O}}$ (3)
where $\text{head}_{i}=\text{Attn}(\textbf{{X}}_{Q}\textbf{{W}}_{i}^{Q},\textbf{{X}}_{K}\textbf{{W}}_{i}^{K},\textbf{{X}}_{V}\textbf{{W}}_{i}^{V})$, with $\textbf{{X}}_{Q},\textbf{{X}}_{K},\textbf{{X}}_{V}\in\mathbb{R}^{N\times d_{k}}$ set to $\textbf{{X}}_{e}$ for self-attention, using trainable weights $\textbf{{W}}_{i}^{Q},\textbf{{W}}_{i}^{K},\textbf{{W}}_{i}^{V}\in\mathbb{R}^{d_{k}\times d_{k}}$ and $\textbf{{W}}^{O}\in\mathbb{R}^{Hd_{k}\times Hd_{k}}$, as depicted in Figure 2.
Figure 2: Architecture of multi-head self-attention.
Next, the MHA output is processed by the FFN, applied uniformly across positions:
$\text{FFN}\left(\textbf{{X}}\right)=\textbf{{W}}_{2}\cdot\text{ReLU}\left(\textbf{{W}}_{1}\cdot\textbf{{X}}+\textbf{{b}}_{1}\right)+\textbf{{b}}_{2}$ (4)
where $\textbf{{W}}_{1}\in\mathbb{R}^{d_{f}\times d_{h}}$, $\textbf{{W}}_{2}\in\mathbb{R}^{d_{h}\times d_{f}}$, $\textbf{{b}}_{1}$, $\textbf{{b}}_{2}$ are trainable parameters, $d_{h}=H\cdot d_{k}$ and $\text{ReLU}(x)=\max(0,x)$ enables complex feature learning, with $d_{f}>d_{h}$ for enhanced performance.
Finally, residual connections and layer normalization link the MHA and FFN layers to produce the output $\textbf{{X}}_{2}$ of a single-layer Transformer encoder:
$\textbf{{X}}_{1}$ $=\text{MHA}\left(\text{LayerNorm}\left(\textbf{{X}}_{e}\right)\right)+\textbf{{X}}_{e}$ (5)
$\textbf{{X}}_{2}$ $=\text{FFN}\left(\text{LayerNorm}\left(\textbf{{X}}_{1}\right)\right)+\textbf{{X}}_{1}$
2.2 Forward Error Correction Codes
In information theory, FEC is a technique designed to reduce the bit error rate (BER) when transmitting data over a noisy channel. We adopt a linear block code defined by a generator matrix $\textbf{G}\in\mathbb{R}^{k\times n}$ and a parity-check matrix $\textbf{H}\in\mathbb{R}^{(n-k)\times n}$, satisfying $\textbf{G}\cdot\textbf{H}^{T}=\textbf{0}$ over $\mathbb{F}_{2}$.
The encoder maps a message vector $\textbf{{m}}\in\{0,1\}^{k}$ to a codeword $\textbf{{x}}\in\{0,1\}^{n}$ via $\textbf{{x}}=\textbf{{m}}\cdot\textbf{G}$, where $\textbf{H}\cdot\textbf{{x}}=\textbf{0}$. This codeword is modulated using Binary Phase Shift Keying (BPSK), defined as $\text{BPSK}(x)=1-2x$, yielding $\textbf{{x}}_{s}$, which is transmitted over Binary-Input Discrete Memoryless Channels (B-DMCs, e.g., AWGN), producing $\textbf{{y}}=\textbf{{x}}_{s}+\textbf{{n}}$, with $\textbf{{n}}\sim\mathcal{N}(0,\sigma^{2})$.
Figure 3: Architecture of the communication system.
Figure 4: Decoder framework with pre- and post-processing.
The decoder aims to recover m from y, as depicted in Figure 3. Our work focuses on designing and training a parameterized decoding function $f_{\theta}$. To develop a model-free Transformer-based decoder, we leverage pre- and post-processing techniques from  [28], preserving BER and Mean Squared Error (MSE) performance (see Figure 4). Training with an all-zero codeword mitigates overfitting due to exponential codeword space growth. And the B-DMCs is modeled by (6) which is an equivalent statistical mode that differs from the true physical one  [29].
$\textbf{{y}}=\textbf{{x}}_{s}\cdot\textbf{{z}}$ (6)
where z is a random multiplicative noise independent of $\textbf{{x}}_{s}$}.
In the pre-processing step, the absolute values of received signal and the syndromes are concatenated as the input to the neural network, which can be written as:
$\widetilde{\textbf{{y}}}=\left[\lvert\textbf{{y}}\rvert,\text{s}\left(\textbf{{y}}\right)\right]$ (7)
where, $\left[\cdot,\cdot\right]$ denotes the concatenation of vectors, $\lvert\textbf{{y}}\rvert$ signifies the absolute value of y and $\text{s}(\textbf{{y}})=\textbf{H}\cdot\text{bin}(\text{sign}(\textbf{{y}}))\in\{0,1\}^{n-k}$ is the syndrome from hard-decoded y.
In the post-processing step, the predicted noise $\hat{\textbf{{z}}}$ is multiplied by the received signal y to recover x. That is, the predicted $\hat{\textbf{{x}}}$ takes the form:
[…]
2.3 Mask Generation in ECCT and FECCT
Masks in Transformer-based error correction enhance decoding by integrating code structure into attention mechanisms. In ECCT  [24], the mask $g(\textbf{H}):\{0,1\}^{(n-k)\times n}\to\{-\infty,0\}^{(2n-k)\times(2n-k)}$ embeds code structure using the parity-check matrix H. Initialized as an identity matrix, it unmasks positions of paired ones in H rows, extending the Tanner graph to two-ring connectivity. It is applied additively in attention: $\text{Attn}(\textbf{{Q}},\textbf{{K}},\textbf{{V}})=\text{Softmax}\left(\frac{\textbf{{Q}}\cdot\textbf{{K}}^{T}+g(\textbf{H})}{\sqrt{d_{k}}}\right)\cdot\textbf{{V}}$.
In FECCT  [25], a learned $\psi:\mathbb{N}\to\mathbb{R}$ modulates attention via the distance matrix $\mathcal{G}(\textbf{H})\in\mathbb{N}^{(2n-k)\times(2n-k)}$, applied with a Hadamard product: $\text{Attn}(\textbf{{Q}},\textbf{{K}},\textbf{{V}})=\left(\text{Softmax}\left(\frac{\textbf{{Q}}\cdot\textbf{{K}}^{T}}{\sqrt{d_{k}}}\right)\odot\psi(\mathcal{G}(\textbf{H}))\right)\cdot\textbf{{V}}$.
3 Proposed Unified Error Correction Code Transformer
In this section, we present the complete architecture and training process of the proposed Unified Error Correction Code Transformer (UECCT).
3.1 Shared Unified Attention
The vanilla self-attention mechanism in Transformer architectures establishes dense correlations among all input features within a sequence, resulting in a computational complexity of $\mathcal{O}(N^{2}\cdot d_{k}+N\cdot d_{k}^{2})$, where $N$ is the sequence length and $d_{k}$ is the embedding dimension. This approach, while effective for capturing long-range dependencies, is computationally intensive and overlooks the intrinsic relationships across different samples, particularly in channel decoding tasks involving diverse linear block codes. Moreover, prior studies [30] suggest that the input to the softmax function in self-attention is probabilistically sparse, with only a few weights dominating the attention distribution. This sparsity implies that a low-rank approximation could effectively capture the dominant correlations, reducing redundancy without compromising performance.
Figure 5: Rank distribution of the $\textbf{{Q}}\cdot\textbf{{K}}^{T}$ matrix in the ECCT model across 6 encoder layers for various code types.
To investigate this hypothesis, we trained an ECCT model [24] with parameters $L=6$ (number of Transformer encoder layers), $H=8$, $d_{k}=\text{max}(2n-k)=182$, and $d_{f}=2912$ on a diverse set of linear block codes, including Polar(32,16), Polar(64,32), Polar(128,86), LDPC(49,24), LDPC(121,60), LDPC(121,80), BCH(31,16), BCH(63,36), and BCH(127,120). We analyzed the rank distribution of the $\textbf{{Q}}\cdot\textbf{{K}}^{T}$ matrix across 1000 batches for the six-layer Transformer encoder. As illustrated in Figure 5, the rank ranges from 12 to 181, with over 90% of values below 69, consistently indicating a low-rank structure across all layers and code types. This observation suggests that the high-dimensional self-attention matrix can be approximated with a lower-rank representation, offering a pathway to reduce computational overhead.
Figure 6: Distribution of Jensen-Shannon divergence between attention head pairs across 6 layers, showing high similarity (JSD << 0.17).
Furthermore, the conventional multi-head self-attention mechanism computes independent attention scores for each head using distinct weight matrices $\textbf{{W}}_{i}^{Q}$, $\textbf{{W}}_{i}^{K}$, and $\textbf{{W}}_{i}^{V}$. However, we hypothesize that the attention distributions across heads may exhibit high similarity, particularly in channel decoding, where the sparse structure of the parity-check matrix $\mathbf{H}$ constrains bit interactions. To quantify this, we employed the Jensen-Shannon Divergence (JSD) [31], a symmetric measure of similarity between probability distributions, defined as:
$\text{JSD}(\mathbf{A}_{i}[k]\|\mathbf{A}_{j}[k])=\frac{1}{2}\left[D_{KL}(\mathbf{A}_{i}[k]\|\mathbf{M}[k])+D_{KL}(\mathbf{A}_{j}[k]\|\mathbf{M}[k])\right]$ (9)
where $k$ is the row index, $\mathbf{A}_{i}$ and $\mathbf{A}_{j}$ $\in\mathbb{R}^{N\times N}$are the softmax outputs of attention heads $i$ and $j$, $\mathbf{M}[k]=\frac{1}{2}(\mathbf{A}_{i}[k]+\mathbf{A}_{j}[k])$ is the average distribution, and $D_{KL}(\mathbf{P}\|\mathbf{Q})=\sum_{l=1}^{N}\mathbf{P}[l]\log\frac{\mathbf{P}[l]}{\mathbf{Q}[l]}$ is the Kullback-Leibler divergence [32]. Using the same simulation setup as the low-rank validation ($L=6$, $H=8$, $d_{k}=182$, $d_{f}=2912$), we computed the per-row JSD for all head pairs across all rows, averaging the results. As shown in Figure 6, JSD values are consistently below 0.17 across all layers, with a mean of 0.13, indicating strong alignment in attention distributions. This similarity, driven by the sparse structure of $\mathbf{H}$, supports the feasibility of sharing a common attention matrix across heads.
Motivated by these findings, we propose a novel Unified Attention Module (UAM) to address the computational inefficiency and lack of cross-code generalization in traditional self-attention. The UAM introduces learnable memory components, $\textbf{{A}}_{l}$ and $\textbf{{V}}_{l}\in\mathbb{R}^{N\times d_{l}}$, where $d_{l}$ is a hyperparameter controlling the low-rank approximation. The UAM is defined as:
$\text{UniAttn}\left({{\textbf{{A}}_{l}},{\textbf{{V}}_{l}}}\right)=\text{Softmax}\left({{\textbf{{A}}_{l}}}\right)\cdot\left({{\textbf{{V}}_{l}^{T}}\textbf{{X}}}\right)$ (10)
where $\textbf{{X}}\in\mathbb{R}^{N\times d_{k}}$ is the input to the attention layer. In contrast, the conventional self-attention is expressed as:
$\text{Attn}\left(\textbf{{W}}^{\textit{Q}},\textbf{{W}}^{\textit{K}},\textbf{{W}}^{\textit{V}}\right)=\text{Softmax}\left(\frac{\left(\textbf{{XW}}^{\textit{Q}}\right)\cdot{\left(\textbf{{XW}}^{\textit{K}}\right)}^{T}}{\sqrt{d_{k}}}\right)\cdot\left(\textbf{{XW}}^{\textit{V}}\right)$ (11)
By eliminating the projections $\textbf{{W}}^{\textit{Q}}$ and $\textbf{{W}}^{\textit{K}}$, the UAM reduces the computational complexity from $\mathcal{O}(N^{2}\cdot d_{k}+N\cdot d_{k}^{2})$ to $\mathcal{O}(N\cdot d_{l}\cdot d_{k})$, where $d_{l}\ll N$. The shared attention matrix $\mathbf{A}_{l}$ is applied across all heads, leveraging the observed head similarity to foster a cohesive decoding architecture. This design not only reduces the parameter count but also enables the model to learn shared features across diverse code types, facilitating a code-agnostic framework.
Figure 7: Architecture of the vanilla multi-head self-attention (A) and the proposed multi-head unified-attention (B).
Algorithm 1 Pseudocode for Multi-Head Unified-Attention
1 Input: X  # $\text{shape}=(B,N,d_{h})$
2 Trainable Parameters: $\textbf{{A}}_{l}$, $\textbf{{V}}_{l}$
3 Hyper Parameters: number of heads $H$; batch size $B$
4 Hyper Parameters: embedding size $d_{k}$; $d_{h}=H\times d_{k}$
5 $\textbf{{X}}=\textbf{{X}}.\text{view}(B,N,H,d_{k}).\text{transpose}(1,2)$
6 $\textbf{{A}}=\text{Softmax}(\textbf{{A}}_{l},\text{dim}=-1)$  # $\text{shape}=(B,1,N,d_{l})$
7 $\textbf{{X}}_{o}=\textbf{{A}}\cdot\left(\textbf{{V}}_{l}^{T}\cdot\textbf{{X}}\right)$  # $\text{shape}=(B,H,N,d_{k})$
8 $\textbf{{X}}_{o}=\textbf{{X}}_{o}.\text{transpose}(1,2).\text{view}(B,N,d_{h})$
9 $\textbf{{X}}_{o}=\textbf{{W}}^{O}(\textbf{{X}}_{o})$
10 Output: $\textbf{{X}}_{o}$  # $\text{shape}=(B,N,d_{h})$
The architecture of the proposed multi-head UAM is depicted in Figure 7, with the corresponding pseudocode provided in Algorithm 1. The shared attention mechanism, combined with the low-rank approximation, significantly reduces model complexity while preserving decoding accuracy, as validated in subsequent experiments. This approach marks a significant advancement over the vanilla Transformer-based ECCT [24], which lacks cross-code unification.
3.2 Sparse Masked Unified Attention
Error detection and correction in channel decoding rely on analyzing received codewords against parity-check equations, where a non-zero syndrome indicates channel errors. In Transformer architectures, vanilla self-attention is inherently dense, correlating all bits indiscriminately, despite parity bits not interacting with every information bit. To address this, we propose a sparse masked unified attention mechanism that incorporates code-specific dependencies, inspired by human intelligence in discerning bit relationships. This approach accelerates the model’s understanding of decoding principles, enhancing performance.
Specifically, for any linear block code defined by a parity-check matrix $\textbf{H}\in\mathbb{F}_{2}^{(n-k)\times n}$, we define a mapping $\text{M}(\textbf{H}):\{0,1\}^{(n-k)\times n}\to\{-\infty,0\}^{N\times d_{l}}$ that generates a mask for the attention memory $\textbf{{A}}_{l}$. The UAM is formulated as:
$\text{UniAttn}\left(\boldsymbol{A}_{l},\boldsymbol{V}_{l}\right)=\text{Softmax}\left(\boldsymbol{A}_{l}+\text{M}(\textbf{H})\right)\cdot\left(\boldsymbol{V}_{l}^{T}\boldsymbol{X}\right)$ (12)
The design of the mask function $\text{M}(\textbf{H})$ aims to enable the neural network to capture the relationship between the received signal $\textbf{{y}}\in\mathbb{R}^{n}$ and its syndrome $\text{s}(\textbf{{y}})=\textbf{H}\cdot\overline{\textbf{{y}}}\in\{0,1\}^{n-k}$, where $\overline{\textbf{{y}}}=\text{bin}(\text{sign}(\textbf{{y}}))$ is the hard-decoded binary vector derived from y. For a valid codeword y (i.e., noise-free or correctly decoded), the orthogonality condition $\textbf{H}\cdot\overline{\textbf{{y}}}=\textbf{0}$ holds over $\mathbb{F}_{2}$, yielding $\text{s}(\textbf{{y}})=\textbf{0}$. To embed this relationship within the attention mechanism, we define an extended parity-check matrix $\overline{\textbf{H}}\in\mathbb{R}^{(2n-k)\times(n-k)}$ satisfying:
$\left[\overline{\textbf{{y}}},\text{s}\left(\textbf{{y}}\right)\right]\cdot\overline{\textbf{H}}=\textbf{0}$ (13)
Inspired by this, we explore a mask structure leveraging the parity-check constraints. Let $\textbf{{A}}_{l}$ be a matrix in $\mathbb{R}^{N\times d_{l}}$, where $N$ is defined as 2nk2n-k and $d_{l}$ as nkn-k. The theoretical foundation for setting $d_{l}\approx n-k$ stems from the maximum rank of the parity-check matrix $\textbf{H}\in\mathbb{F}_{2}^{(n-k)\times n}$, which is nkn-k. We then construct the extended parity-check matrix $\overline{\textbf{H}}$ as follows:
$\overline{\textbf{H}}=\begin{bmatrix}{\textbf{H}}^{T}\\ {\textbf{{I}}}_{n-k}\end{bmatrix}$ (14)
where ${\textbf{{I}}}_{n-k}\in\mathbb{R}^{(n-k)\times(n-k)}$ is an identity matrix with ones on the diagonal and zeros elsewhere. The generation of the attention mask is thus a direct process, formulated as $\text{M}(\overline{\textbf{H}})$.
We naturally proceed to define the density of $\overline{\textbf{H}}$ as:
$H_{d}=\frac{\text{sum}(\overline{\textbf{H}})}{(2n-k)\times(n-k)}$ (15)
Figure 8: We present the corresponding matrices $\textbf{H}^{T}$, $\overline{\textbf{H}}$ and the Tanner graph for the Hamming (7,4) code.
For enhanced clarity, Figure 8 offers a visual depiction of matrices $\textbf{H}^{T}$, $\overline{\textbf{H}}$ and the Tanner graph, utilizing the Hamming (7,4) code as an illustrative example. Here, matrix H is defined as:
$\textbf{H}=\begin{bmatrix}1&1&1&0&1&0&0\\ 1&0&1&1&0&1&0\\ 0&1&1&1&0&0&1\end{bmatrix}$ (16)
Algorithm 2 Pseudocode for Sparse Mask Construction
1 Input: H  # $\text{shape}=(n-k,n)$
2 $\overline{\textbf{H}}=\text{zeros}(2n-k,n-k)$
3 $\overline{\textbf{H}}[0:n,0:n-k]=\textbf{H}.\text{transpose}(0,1)$
4 $\overline{\textbf{H}}[n:2n-k,0:n-k]=\text{eye}(n-k)$
5 Output: $(-\infty)\cdot(\neg\mathrm{~\overline{\textbf{H}}})$
This illustration demonstrates that the application of the mask results in sparser attention, a phenomenon particularly pronounced in low-density parity-check codes. Most importantly, the complexity of the UAM is now reduced to the density of the code $\mathcal{O}(N\times d_{l}\times d_{k}\times H_{d})$. Our experimental results indicate that the use of this mask has successfully reduced the computational complexity of the attention mechanism by an average of 86%. Furthermore, we provide a summary of the Python-style sparse mask construction within Algorithm 2. It should be noted that, in Application-Specific Integrated Circuit (ASIC) implementations, computations involving the masked parts can be directly skipped to enhance efficiency.
3.3 Architecture and Training Methodology
The architecture of the proposed unified error correction code Transformer is illustrated in Figure 9. The primary innovation in our model involves substituting the vanilla attention mechanism with a multi-head shared unified-attention module. This modification not only decreases computational complexity but also empowers the model to discern distinctive features among diverse codewords, thereby facilitating the establishment of a unified decoding architecture. We employ fine-tuning techniques to enhance the model’s generalization to unseen codewords. Furthermore, we have designed standardization units to standardize the lengths of codewords and syndromes across diverse datasets, ensuring the model’s compatibility with various code types, lengths, and rates.
Figure 9: Illustration of the proposed unified error correction code Transformer.
In Figure 9, the output y from the AWGN channel passes through the pre-processing module to yield N=2nkN=2n-k input elements for the neural network. The decoder is constructed by concatenating $L$ Transformer encoders, which consist of unified-attention and feed-forward layers, with normalization layers in between. The output module is characterized by two fully connected layers. The initial layer compresses the multi-head output into a one-dimensional vector of size 2nk2n-k, with the subsequent layer transforming it into an $n$-dimensional vector representing the softly decoded noise. Finally, the post-processing module acts on the estimated noise to derive the transmitted original codeword.
We employ binary cross-entropy as the loss function with the objective of learning to predict the multiplicative noise z. The corresponding target binary multiplicative noise is denoted as $\tilde{\textbf{{z}}}=\text{bin}(\text{sign}(\textbf{{z}}))$. Therefore, the loss calculation for a received individual codeword y is given by:
$\text{loss}=-\sum_{i=1}^{n}{\tilde{z}}_{i}\text{log}(f_{\theta}(y))+(1-{\tilde{z}}_{i})\text{log}(1-f_{\theta}(y))$ (17)
To accommodate codewords of varying lengths within a unified architecture, we have designed a standardization unit that performs padding on $\lvert\textbf{{y}}\rvert$ and $s(\textbf{{y}})$ as described in Equation (7). Specifically, after receiving the channel output, we pad different codewords and their corresponding syndromes with zeros up to the maximum length. Assuming that the maximum length of the codeword is $N_{\text{max}}$ and the maximum length of the syndrome is $S_{\text{max}}$, the dataset is standardized according to:
$\widetilde{\textbf{{y}}}=\left[~\left[\lvert\textbf{{y}}\rvert,~\textbf{0}_{c}\right],~\left[s(\textbf{{y}}),~\textbf{0}_{s}\right]~\right]$ (18)
where $\textbf{0}_{c}$ and $\textbf{0}_{s}$ represent zero vectors with sizes $N_{\text{max}}-\text{size}(\textbf{{y}})$ and $S_{\text{max}}-\text{size}(s(\textbf{{y}}))$, respectively. Through this unit, the Transformer neural decoder can accommodate any codeword within the maximum length, allowing for mixed training in a unified manner. In ASIC implementations, computations for zero-padded parts can be skipped to improve efficiency, similar to the sparse-masked attention.
In the training phase, we conducted 1000 epochs, each consisting of 1000 minibatches, which in turn contained 512 samples each. Samples for each minibatch were randomly selected from a pool of systematic Polar [33], LDPC, and BCH codewords. The learning rate was initialized at $10^{-3}$, and a cosine decay scheduler was applied without warmup  [34], gradually reducing it to $10^{-6}$ by the end of the training process. We employ the Adam optimizer to dynamically adjust the learning rate, leveraging its efficacy in managing diverse data types and model architectures  [35]. The AWGN noise introduced to the links had its Signal-to-Noise Ratio (SNR) randomly chosen between 3 and 7 for each batch. All experiments were conducted on NVIDIA GeForce RTX 4090 24GB GPUs.
4 Experiments
We conducted experiments on linear block codes, specifically Polar, LDPC, and BCH, to assess the effectiveness of the proposed architecture. The parity-check matrices utilized in the paper were sourced from  [36]. We benchmark our approach against the most recent state-of-the-art performance from  [24], which trains each codeword parameter independently, and the hybrid-trained FECCT  [25]. Additionally, we compare it to legacy BP-based decoders, including unlearned BP  [37], fully supervised Hyper BP  [38], and Autoregressive BP (AR BP)  [39]. The results are presented in the form of bit error rate (BER) and block error rate (BLER) at various normalized SNR (i.e. $E_{b}/N_{0}$) values, with at least $10^{5}$ random codewords tested per code type. We define aggregate BER (ABER) and aggregate BLER (ABLER) as the overall error rates spanning all evaluated codewords.
In addition, to enable direct comparison with ECCT and FECCT, whose simulation results were directly obtained from their respective papers  [24, 25], we fine-tuned the pre-trained UECCT models on Consultative Committee for Space Data Systems (CCSDS)  [40] and MacKay  [41] codewords leveraging the hybrid-trained Polar, LDPC, and BCH models. Given GPU memory constraints and the fact that the performance of long codes has been thoroughly explored  [42], while the error correction of short codes remains far from the Shannon limit  [43], this work primarily focuses on training and validating medium-to-short codes.
The proposed unified error correction code Transformer is defined by the following hyperparameters: the number of Transformer encoder layers $L$, the number of unified attention heads $H$, the dimensions of word embeddings $d_{k}$, the dimensions $d_{l}$ of the trainable memory $\textbf{{A}}_{l}$ and $\textbf{{V}}_{l}$, and the dimension $d_{f}$ of the FFN layer. We present the performance of our framework with hyperparameters are set to $L=6$, $H=8$, $d_{k}=64$, $d_{l}=64$, and $d_{f}=4\cdot H\cdot d_{k}$. The value of $d_{l}=64$ corresponds to the maximum syndrome length among the trained codewords.
Figure 10: BER comparison of ECCT, FECCT, and proposed UECCT for Polar codes.
Figure 11: BER comparison of ECCT and proposed UECCT for LDPC codes.
Figure 12: BER comparison of ECCT, FECCT, and proposed UECCT for BCH codes.
Table 1: A comparison of the negative natural logarithm of the BER
Table 1 presents the simulation results, specifically showing the negative natural logarithm of the BER at $E_{b}/N_{0}$ values of 4, 5, and 6 dB. It is noted that higher values in this context indicate better performance, with the best results for each SNR highlighted in bold. Additional BER plots for Polar, LDPC and BCH codes are provided in Figure 10, 11 and Figure 12, respectively. Simulation results demonstrate that our proposed UECCT outperforms both the individually trained ECCT  [24] and the hybrid-trained FECCT  [25] across the majority of codewords, validating the effectiveness of the proposed architecture. Notably, UECCT even surpasses FECCT’s performance with an embedding length of 128 when operating at a more compact embedding length of 64, further highlighting the efficiency and power of our design. The superior performance stems from its ability to leverage a shared attention mechanism across heads, as detailed in Section 3, combined with sparsity-aware masking that reduces computational complexity while preserving decoding accuracy.
5 Analysis
We investigate the impact of Transformer encoder layers on performance, focusing on the embedding dimension $d_{k}$ and the effectiveness of sparse masked attention. Additionally, we compare the computational complexity of our approach with existing methods.
5.1 Impact of Encoder Layers and Embedding Dimensions
Figure 13 displays the ABER performance across varying encoder layers $L$ and embedding dimensions $d_{k}$. Results reveal that $L$ exerts a stronger influence on performance than $d_{k}$. Increasing $L$ broadens the parameter space, enhancing nonlinear fitting and enabling complex representations.
Figure 13: ABER performance across various Transformer encoder layers and the embedding dimensions $d_{k}$.
Deeper architectures better capture intricate data patterns and dependencies, crucial for error correction. Additional layers enable multi-level abstraction and refinement, improving error interpretation and correction, especially in complex wireless communication systems where simpler models falter.
These findings highlight the primacy of model depth in Transformer-based error correction. Prioritizing $L$ over $d_{k}$ can optimize model design, enhancing practical performance.
5.2 Impact of Sparse Masked Attention
Figure 14 illustrates the effect of sparse masked attention on training loss, ABER, and ABLER. Sparse masked attention reduces training loss by 33.3% and speeds up convergence. It also significantly boosts error correction, improving ABER by 0.7 dB at a bit error rate of $10^{-5}$ and ABLER by 0.9 dB at a block error rate of $10^{-4}$.
Figure 14: The overall improvement in loss, ABER and ABLER with sparse masked attention.
These gains highlight the effectiveness of embedding domain-specific knowledge into the attention mechanism, yielding significant improvements in efficiency and accuracy. By reducing computational complexity while boosting the model’s capacity to decode complex signals, this approach emerges as a promising advancement for future error correction techniques.
5.3 Computational Complexity Analysis
Before delving into the computational complexity, let’s first review the unified error correction code Transformer presented in Figure 9, which incorporates $L$ layers of Transformer encoders. The principal computational complexity within each layer is found in the multi-head self-attention module, which is predominantly responsible for matrix multiplication and is defined by the hyperparameters $H$, $d_{k}$ and $d_{l}$. Consequently, the computational complexity of the proposed architecture is given by $\mathcal{O}\left(L\times H\times\left(N\times d_{l}\times d_{k}\times H_{d}\right)\right)$, where N=2nkN=2n-k, and $d_{l}\ll N$. Thanks to the sparsity of the parity-check matrix, $H_{d}$ is usually quite small. In our experiment, $H_{d}$ is approximately 0.14, which significantly reduces the computational complexity of the neural network. By contrast, the vanilla Transformer architectures in ECCT and FECCT have complexities of $\mathcal{O}\left(L\times H\times\left(N^{2}\times d_{k}\times M_{d}+N\times d_{k}^{2}\right)\right)$ and $\mathcal{O}\left(L\times H\times\left(N^{2}\times d_{k}+N\times d_{k}^{2}+N^{2}\times M_{d}\right)\right)$, respectively, where $M_{d}$ represents the mask density in these models.
Table 2: Comparison of trainable parameters, MACs, training time, inference time, and mask density among ECCT, FECCT, and UECCT for various codes, all with identical parameters $L=6$, $H=8$, $d_{k}=64$, $d_{l}=64$, and $d_{f}=2048$
Table 2 compares the Trainable Parameters (TPs), Multiply-Accumulate Operations (MACs), training time, inference time, and mask density for ECCT, FECCT, and UECCT. Lower values are preferable, with the best performing values highlighted in bold. The metrics for TPs and MACs are obtained using Thop  [44], a neural network profiling tool. UECCT reduces parameters and MACs by 25.0% compared to both ECCT and FECCT through low-rank approximation, shared attention, and omission of Q and K projections. This results in training times 19.7% and 81.2% lower, and inference times 23.1% and 93.5% lower than ECCT and FECCT, respectively. Mask densities are influenced by design choices: FECCT has the highest density due to Tanner graph distance calculations, ECCT has a moderate density from extended bit-pair relations, and UECCT has the lowest density, 58.2% and 79.7% lower than ECCT and FECCT, achieved through the direct utilization of H and the extension of the identity matrix. For a more intuitive representation, the average values of each metric across the codewords in Table 2 are provided in Figure 15.
Figure 15: Comparison of average TPs, MACs, training time, inference time, and mask density between ECCT, FECCT and UECCT.
Compared to the approach in  [24, 25], while the proposed method excels in decoding versatility and computational complexity, it may fall short in terms of memory efficiency, power consumption, and computational resource utilization when compared to non-learning solutions. This inefficiency could limit its potential for deployment. To address these issues, future implementations could leverage techniques such as parameter sharing, low-rank factorization, and quantization, aiming to reduce complexity and enhance efficiency  [45, 46].
6 Conclusions
We introduced a unified, code-agnostic Transformer decoding architecture, designed to integrate various linear block codes seamlessly and offer superior decoding performance. Our approach leverages the self-attention mechanism inherent in Transformer models, enabling effective handling of sequential data and capturing long-range dependencies critical for decoding tasks. Through rigorous analysis and extensive experimental validation, our results indicate that the unified Transformer decoding architecture can significantly advance the state-of-the-art in decoding performance, potentially setting new benchmarks for future research and practical applications.
In conclusion, the unified Transformer decoding architecture presented in this paper signifies a significant step forward in the field of error correction coding. It paves the way for the development of more adaptable, efficient, and robust decoding solutions, poised to meet the stringent requirements of 6G and beyond. As the field progresses, we anticipate that the continued integration of AI technologies will unlock new possibilities, driving innovation in wireless communication systems. Future work can aim to further reduce hardware implementation complexity and enhance computational efficiency using techniques such as quantization and compression of neural network models.